OpenStudy (anonymous):

Gail paid $118.65 for a varsity jacket. The price includes a 5% sales tax. How much was the jacket before the tax? a. $112.72 b. $113.00 c. $117.00 d. $124.65 I need help.

OpenStudy (anonymous):

Let x be the cost before tax. From the text you get the information that x + 0,05*x = $ 118,65

OpenStudy (anonymous):

Nevermind, I think I got the answer, but thanks.

OpenStudy (anonymous):

Did you divide 118,65 by 1,05?

OpenStudy (anonymous):

No, I actually found the answer by : what is 5% of 118.65. I set up the problem, I got 5.93 is 5% of $118.65 so I subtracted $5.93 from 118.65 and got $112.72

OpenStudy (hossamhoussien):

\[\frac{ 5 }{ 100 } \times118.65=5.9325$\] Then the total amount - tax amount = jacket price \[118.65-5.9325=112.7175\approx112.72\]:

OpenStudy (anonymous):

You guys are doing it incorrect... Here's an example why that doesn't work: Guy paid $105 for a varsity jacket. The price includes a 5% sales tax. How much was the jacket before the tax? Your method: 0,05 * 105 = 5,25 105 - 5,25 = 99,75 My method: x + 0,05x = 105 1,05x = 105 x = 105/1,05 = 100

OpenStudy (anonymous):

Thank You all!